# Measuring Social Biases in Masked Language Models by Proxy of Prediction Quality

### v2

@@ Abstract
Innovative transformer-based language models produce contextually-aware token embeddings and have achieved state-of-the-art performance for a variety of natural language tasks, but have been shown to encode unwanted biases for downstream applications. In this paper, we evaluate the social biases encoded by transformers trained with the masked language modeling objective using proposed proxy functions within an iterative masking experiment to measure the quality of transformer models’ predictions, and assess the preference of MLMs towards disadvantaged and advantaged groups. We compare bias estimations with those produced by other evaluation methods using benchmark datasets and assess their alignment with human annotated biases. We find relatively high religious and disability biases across considered MLMs and low gender bias in one dataset relative to another. We extend on previous work by evaluating social biases introduced after retraining an MLM under the masked language modeling objective, and find that proposed measures produce more accurate estimations of biases introduced by retraining MLMs than others based on relative preference for biased sentences between models.
@@ 1 Introduction
Masked language models (MLM(s); Brown et al., 2020; Devlin et al., 2019; Liu et al., 2019a; Radford et al., 2019) such as transformers BERT and RoBERTa incorporate bidirectionality and self-attention, producing contextually-aware embeddings. Unlike $\text{BERT}_{\text{unc}}$, RoBERTa (Robustly Optimized BERT) was pretrained solely on the masked language modeling objective (MLMO) on a larger corpus of text. RoBERTa uses a dynamic masking strategy for a diverse set of representations during training and has achieved state-of-the-art results on GLUE, RACE, and SQuAD (Liu et al., 2019b; Rajpurkar et al., 2016; Lai et al., 2017; Wang et al., 2018). Distilled model variants have been shown to train significantly faster for minor decreases in performance Sanh et al. (2019). MLMs have produced state-of-the-art results for masked language modeling, named entity recognition, and intent or topic classification, but also encode concerning social biases against disadvantaged groups that are undesirable in production settings. As MLMs become increasingly prevalent, researchers have been working on methods to measure biases embedded in these models Nangia et al. (2020); Kaneko and Bollegala (2022); Salutari et al. (2023).
@@
To address the issue of social biases in MLMs at its source, we measure biases of MLMs while focusing on an MLM’s key pretraining objective, masked language modeling. In this work, we focus on proposing and assessing bias evaluation measures for MLMs and not proposing methods for debiasing MLMs. However, when assessing evaluation measures, we consider prior research that involves retraining or fine-tuning under the MLMO with debiased or counterfactual data to reduce biases in MLMs, such as Zhao et al., 2018 and Zhao et al., 2019 which use data augmentation to swap gendered words with their opposites prior to retraining. Motivated by this, we assess whether proposed measures satisfy an important criterion for improvement over previously proposed ones: alignment with biases introduced by MLM retraining under the MLMO (Section 3.5.2).
@@
We represent MLM bias through a model’s relative preference for ground truth tokens between two paired bias sentences, where one contains bias against disadvantaged groups (stereotypical) and the other contains bias against advantaged groups (anti-stereotypical). We assess the relative preferences of the MLMs towards sentences biased against disadvantaged and advantaged groups using proxy measures for prediction quality. In particular, we quantify MLM preference by proxy of masked token prediction quality given unmasked token context between encoded sentences within pairs.
@@
To measure the preference of an MLM using the (attention-weighted) quality of its predictions under the MLMO, we propose and validate proxy functions that measure the likelihood an MLM will select a ground truth token to replace a masked one, such as crra (Equation 6) and $\Delta\textsc{pa}$ (Equation 7), and extend these definitions for a sentence. We apply per-model indicator function bspt (Equation 9) to estimate the encoded social biases in pretrained MLMs. Our approach differs from prior research in measuring biases in MLMs by using attention-weights under an iterative masking experiment (ime; Section 3) to probe MLM preferences.
@@
We compare pre- and retrained MLMs within the same model class to recover the nature of biases introduced by retraining. In particular, we define and apply a proxy for the relative preference between two MLMs with model-comparative indicator function bsrt (Bias Score for MLM Re-training; Equation 10) to estimate social biases introduced by retraining MLMs under the MLMO. It is important to acknowledge that these "introduced" biases must be represented by the relative change in biases after retraining an MLM under the MLMO.
@@
  • We propose model-comparative indicator function bsrt to estimate the social biases against disadvantaged groups for a retrained MLM relative to its pretrained base, and assess bias evaluation measures for alignment with biases introduced by MLM retraining under the MLMO.
  • We evaluate social biases for four transformer models available through the Transformers library Wolf et al. (2020). We use proxy measures for MLM prediction quality with model-comparative function bsrt to estimate social biases introduced by retraining MLMs under the MLMO. We find that proposed measures produce more accurate estimations of biases introduced by retraining MLMs than previously proposed ones, which can produce concerning underestimations of biases after retraining MLMs on sentences biased against disadvantaged groups.
Our methodology could help others evaluate social biases encoded in an MLM after it is retrained on the MLMO, such as for any downstream fill-mask task. We release a package for measuring biases in MLMs to facilitate bias score computation on user-supplied or benchmark datasets and enable easy integration into existing evaluation pipelines for MLMs.
@@ 2 Related Work
2.1 Biases in Non-contextual Word Embeddings
Non-contextual word (token) embeddings Pennington et al. (2014); Mikolov et al. (2013), can be shifted in a direction to decompose bias embedded in learned text data representations. These could be analogies or biases along an axis, such as gender (Bolukbasi et al., 2016) or race (Manzini et al., 2019). The Word Embedding Association Test (WEAT; Caliskan et al., 2017) measures association between targets and attributes using cosine similarity between static word embeddings, but has been shown to overestimate biases by Ethayarajh et al., 2019 which proposed the more robust relational inner product association (RIPA) method, derived from the subspace projection method to debias vectors in Bolukbasi et al., 2016.
@@ 2.2 Evaluating Biases in MLMs
To measure the bias in masked language models, Salutari et al., 2023 tests MLMs in an iterative fill mask setting and proposes the Complementary Reciprocal Rank (crr; Equation 2) for a masked token given its context, using the measure for a sentence (Appendix E; average of all single masked token crrs with token ordering preserved) as a proxy for an MLM’s prediction quality or preference.
Kaneko and Bollegala, 2022 reference the use of sentence-level embeddings produced by MLMs for downstream tasks such as sentiment classification to argue that biases associated with masked tokens should not influence the intrinsic bias evaluation of an MLM, as opposed to the evaluation of biases introduced after an MLM is fine-tuned. They propose evaluation metrics All Unmasked Likelihood (aul; Appendix F) and aul with Attention weights (aula; Appendix G), where aul and aula are generated by requiring the MLM to predict all tokens (unmasked input) to eliminate biases associated with masked tokens under previously proposed pseudo-likelihood-based scoring methods (Nadeem et al., 2021, Nangia et al., 2020), which assumed that masked tokens are statistically independent.
In contrast with Kaneko and Bollegala, 2022, we focus on an MLM’s key pretraining objective, masked language modeling, to measure social biases of the MLM. [^1] Masked language modeling was a pretraining objective for all transformers considered in this work. Next sentence prediction was not used for RoBERTa and similar variants due to relatively lower performance with its inclusion Liu et al. (2019b). In addition, we measure relative changes in biases w.r.t. the intrinsic biases of the same base MLM after retraining under the MLMO (as opposed to fine-tuning). Thus, we argue that biases associated with masked tokens are not undesirable in our case.
2.3 MLM Preference by Prediction Quality
@@
Salazar et al., 2020 uses pseudo log-likelihood scoring to approximate $P(U|M,\theta)$, or the probability of unmodified tokens conditioned on modified ones. Similarly, Nangia et al., 2020 reports CrowS-Pairs Scores (csps; Appendix B), a pseudo-log-likelihood score for an MLM selecting unmodified tokens given modified ones. Nadeem et al., 2021 reports a StereoSet Score (sss; Appendix C), a pseudo-log-likelihood score for an MLM selecting modified tokens given unmodified ones.
Salutari et al., 2023 tests MLMs in an iterative fill mask setting where the model outputs a set of tokens (or the (log)softmax of model logits mapped to tokens) to fill the masked one, starting with the token of highest probability $P(t_{p}|c)$ and, as such, first rank $\rho(t_{p}|c)=1$ in the set of model token predictions (which is limited according to the MLM’s embedding space).
@@
$\begin{split}\textsc{crr}(t|s_{\setminus t_{m}};\theta)&=\bigl{(}\rho(t_{p}|s_{\setminus t_{m}};\theta)^{-1}\\ &\hskip 12.0pt-\rho(t_{m}|s_{\setminus t_{m}};\theta)^{-1}\bigr{)}\\ &=1-\rho(t_{m}|s_{\setminus t_{m}};\theta)^{-1}\\ &=\textsc{crr}(w;\theta)\end{split}$ (2)
$\textsc{crr}(t|s_{\setminus t_{m}};\theta)$ (Equation 2) is another metric for measuring the bias of an MLM for a sentence $s$, where $\rho(t_{p}|s_{\setminus t_{m}})^{-1}$ is the reciprocal rank of the predicted token (and always equal to 1) and $\rho(t_{m}|s_{\setminus t_{m}})^{-1}$ is the reciprocal rank of the masked token. Thus, $\rho(t_{m}|s_{\setminus t_{m}})^{-1}$ provides a likelihood measure for $t_{m}$ being chosen by the model as a candidate token to replace the ground truth (masked) one.
Salutari et al., 2023 defines $\Delta\textsc{p}(s)$ (Appendix D) as the probability difference for a sentence $s$ and $\textsc{crr}(s)$ as the complementary reciprocal rank for a sentence $s$, and claims that metrics based on crr for a sentence $s$ are necessary to fully capture the biases embedded in MLMs. Kaneko and Bollegala, 2022 propose evaluation metrics aul and aula, generating them by predicting all of the tokens in a given unmasked input sentence $s$ (Section 2). By requiring the MLM to simultaneously predict all of the unmasked tokens in a sentence, the researchers aim to avoid selectional biases from masking a subset of input tokens, such as high frequency words (which are masked more often during training).
@@
In contrast with previous methods such as sss and csps, we refrain from using strictly modified or unmodified subsets of input tokens as context, and instead provide all tokens but the ground truth one as context for MLM prediction under each iteration of our masking experiment. In this sense, and similar to the aul measure proposed by Kaneko and Bollegala, 2022, our non-attention-weighted measures consider all tokens equally. Thus, our measures might also benefit from considering the weight of MLM attention as a proxy for token importance when probing for MLM preferences for ground truth tokens between two paired sentences along a social bias axis.
@@
Existing benchmark datasets such as CPS are limited to one ground truth per masked token, so an important consideration is an MLM’s ability to predict multiple plausible tokens for a context that could qualify for concerning social bias but goes unrecognized during evaluation using previously proposed measures. crr could perform better than pseudo-(log)likelihood-based measures for this sensitivity that yields larger relative differences in $\Delta\text{P}(t|c)$ as opposed to $\textsc{crr}(t|c)$, [^2] The lowest rank of any possible ground truth token within a model’s vocabulary is equal to the size of the vocabulary ($1\leq\rho(t_{m}|c)\leq V$ and $0\leq\textsc{crr}(s)\leq 1-(V)^{-1}$ for an MLM with embedding vocabulary of size $V$)., and is deemed critical for evaluation by Salutari et al., 2023 due to the possible uniformity of probabilities generated by a particular MLM w.r.t. others.
3 Experiments and Findings
3.1 Benchmark Datasets for Social Bias
The Crowdsourced Stereotype Pairs Benchmark (CPS; Nangia et al., 2020) contains biased sentences towards historically advantaged and disadvantaged groups along nine forms of social biases. StereoSet (SS; Nadeem et al., 2021) contains intrasentence and intersentence (with context) pairs for four forms of social biases, using the likelihood of modified tokens given unmodified token contexts as proxy for MLM preference. Similarly, CPS contains characteristic words that distinguish sentences within pairs and define the nature of a particular bias towards either advantaged or disadvantaged groups, but instead uses the relative likelihood of unmodified tokens being chosen by the MLM given a modified context (characteristic word) across a sentence pair.
CPS and SS contain biased sentences towards advantaged and disadvantaged groups (see Appendix A for specific bias categories and sentence counts). To probe for biases of interest that are encoded in considered MLMs, the scope of our experiments include all bias categories and sentences pairs in CPS and intrasentence pairs in SS, since intersentence pairs are not masked for bias evaluation (Kaneko and Bollegala, 2022). We estimate the preference of an MLM towards a stereotypical sentence over a less stereotypical one for each bias category in CPS and SS and report corresponding results.
3.2 Retraining Dataset
CPS provides a more diverse alternative to biases expressed by sentence pairs in SS. Biases widely acknowledged in the United States are well represented in CPS [^3] CPS categories are a ”narrowed” version of the US Equal Employment Opportunities Commission’s list of protected categories Nangia et al. (2020). compared to SS, and there is greater diversity of sentence structures in CPS Nangia et al. (2020). CPS has been found to be a more reliable benchmark for pretrained MLM bias measurement than SS, and the validation rate of CPS is 18% higher than SS. Nangia et al. (2020). Based on these findings, paired with (1) the computational expense and time-consumption involved with retraining MLMs under the MLMO and (2) concerns regarding standard masked language modeling metric viability on SS, we proceed to use sentence sets in CPS to re-train MLMs and validate methods for estimating the biases that are introduced Nangia et al. (2020). The data profile for (cased and uncased) sentence sets used to re-train MLMs is shown in Appendix A.2. We re-train MLMs on $\forall s\in S^{\text{dis}}$ or $\forall s\in S^{\text{adv}}$, where $s$ is a sentence biased towards either advantaged ($S^{\text{adv}}$) or disadvantaged groups ($S^{\text{dis}}$), and compare the accuracy of our proposed measures with others.
3.3 Transformer-based Language Models
We denote re- and pretrained transformers as $T_{R}$ and $T_{P}$ respectively. The subscript ${}_{\text{unc}}$ denotes an uncased model. We report and compare results from the following transformer-based language models available through the HuggingFace library Wolf et al. (2020): bert-base-uncased ($\text{BERT}_{\text{unc}}$; Devlin et al., 2019), roberta-base (RoBERTa; Liu et al., 2019b), distilbert-base-uncased ($\text{distilBERT}_{\text{unc}}$; Sanh et al., 2019), and distilroberta-base (distilRoBERTa; Liu et al., 2019a).
3.4 Recovering Social Biases in Pretrained MLMs

### v3

@@ Abstract
Transformer language models have achieved state-of-the-art performance for a variety of natural language tasks but have been shown to encode unwanted biases. We evaluate the social biases encoded by transformers trained with the masked language modeling objective using proposed proxy functions within an iterative masking experiment to measure the quality of transformer models’ predictions and assess the preference of MLMs towards disadvantaged and advantaged groups. We find all models encode concerning social biases. We compare bias estimations with those produced by other evaluation methods using benchmark datasets and assess their alignment with human annotated biases. We extend previous work by evaluating social biases introduced after retraining an MLM under the masked language modeling objective and find proposed measures produce more accurate and sensitive estimations of biases introduced by retraining MLMs based on relative preference for biased sentences between models, while other methods tend to underestimate biases after retraining on sentences biased towards disadvantaged groups.
@@ 1 Introduction
Masked language models (MLM(s); Brown et al., 2020; Devlin et al., 2019; Liu et al., 2019a; Radford et al., 2019) such as transformers BERT and RoBERTa incorporate bidirectionality and self-attention, producing contextually-aware embeddings. Unlike $\text{BERT}_{\text{unc}}$, RoBERTa was pretrained solely on the masked language modeling objective (MLMO) on a larger corpus of text. RoBERTa uses a dynamic masking strategy for a diverse set of representations during training and has achieved state-of-the-art results on GLUE, RACE, and SQuAD (Liu et al., 2019b; Rajpurkar et al., 2016; Lai et al., 2017; Wang et al., 2018). Distilled model variants have been shown to train significantly faster for minor decreases in performance Sanh et al. (2019). MLMs have produced state-of-the-art results for masked language modeling, named entity recognition, and intent or topic classification, but also encode concerning social biases against disadvantaged groups that are undesirable in production settings. As MLMs become increasingly prevalent, researchers have been working on methods to measure biases embedded in these models Nangia et al. (2020); Kaneko and Bollegala (2022); Salutari et al. (2023).
@@
To address the issue of social biases in MLMs at its source, we measure biases of MLMs while focusing on an MLM’s key pretraining objective: masked language modeling. In this work, we focus on proposing and assessing bias evaluation measures for MLMs and not proposing methods for debiasing MLMs. However, when assessing evaluation measures, we consider prior research that involves retraining or fine-tuning under the MLMO with debiased or counterfactual data to reduce biases in MLMs, such as Zhao et al., 2018 and Zhao et al., 2019 which use data augmentation to swap gendered words with their opposites prior to retraining. Motivated by this, we assess whether proposed measures satisfy an important criterion for improvement over previously proposed ones: alignment with biases introduced by MLM retraining under the MLMO (Section 4.5.1).
@@
We represent MLM bias through a model’s relative preference for ground truth tokens between two paired biased sentences, where one contains bias against disadvantaged groups (stereotypical) and the other contains bias against advantaged groups (anti-stereotypical). We assess the relative preferences of MLMs towards sentences biased against disadvantaged and advantaged groups using proxy measures for prediction quality. In particular, we quantify MLM preference by proxy of masked token prediction quality given unmasked token context between encoded sentences within pairs.
@@
To measure the preference of an MLM using the (attention-weighted) quality of its predictions under the MLMO, we propose and validate proxy functions $\Delta\textsc{pa}$ (Equation 7; Probability Difference with Attention) and crra (Equation 6; Complementary Reciprocal Rank with Attention) and introduce a modified $\Delta\textsc{p}$ (Equation 3) to measure the likelihood an MLM will select a ground truth token to replace a masked one, and extend these definitions for a sentence. We apply per-model indicator function bspt (Equation 9) to estimate the encoded social biases in pretrained MLMs. Our approach differs from prior research in measuring biases in MLMs by using attention-weights under an Iterative Masking Experiment (ime; Section 4) to probe MLM preferences.
@@
We compare pre- and retrained MLMs within the same model class to recover the nature of biases introduced by retraining. In particular, we define and apply a proxy for the relative preference between two MLMs with model-comparative indicator function bsrt (Bias Score for MLM Re-training; Equation 10) to estimate social biases introduced by retraining MLMs under the MLMO. [^1] We acknowledge that ”introduced” biases must be represented by the relative change in biases after retraining an MLM under the MLMO.
@@
  • We propose model-comparative indicator function bsrt to estimate social biases against disadvantaged groups for a retrained MLM relative to its pretrained base, and assess bias evaluation measures for alignment with biases introduced by MLM retraining under the MLMO.
  • We evaluate social biases for four transformer models available through the Transformers library Wolf et al. (2020). We use proxy measures for MLM prediction quality with model-comparative function bsrt to estimate social biases introduced by retraining MLMs under the MLMO. We find proposed measures $\Delta\textsc{pa}$ and crra, along with modified $\Delta\textsc{p}$ and crr, produce more accurate estimations of biases introduced by MLM retraining than previously proposed ones, which can produce concerning underestimations of biases after retraining MLMs on sentences biased against disadvantaged groups. We observe that proposed measures $\Delta\textsc{pa}$ and crra, along with modified $\Delta\textsc{p}$, show greater sensitivity than crr, csps, aul, and aula to relative changes in MLM bias due to retraining, indicated by larger and smaller bias scores and more frequently significant relative difference in proportions of bias between pre- and retrained transformers.
Our methodology could help others evaluate social biases encoded in an MLM after it is retrained on the MLMO, such as for any downstream fill-mask task, or after it is fine-tuned for other objectives that alter weights and change MLMO performance. We release a package for measuring biases in MLMs to enable bias score computation on user-supplied or benchmark datasets and easy integration into existing evaluation pipelines.
@@ 2 Related Work
2.1 Biases in Static Word Embeddings
Static word embeddings Pennington et al. (2014); Mikolov et al. (2013) can be shifted in a direction to decompose bias embedded in learned text data representations. These could be analogies or biases along an axis, such as gender (Bolukbasi et al., 2016) or race (Manzini et al., 2019). The Word Embedding Association Test (WEAT; Caliskan et al., 2017) measures association between targets and attributes using cosine similarity between static word embeddings, but has been shown to overestimate biases by Ethayarajh et al., 2019 which proposed the robust relational inner product association (RIPA) method, derived from the subspace projection method to debias vectors in Bolukbasi et al., 2016.
@@ 2.2 Evaluating Biases in MLMs
Salazar et al., 2020 uses pseudo-log-likelihood scoring to approximate $P(U|M,\theta)$, or the probability of unmodified tokens conditioned on modified ones. Similarly, Nangia et al., 2020 reports CrowS-Pairs Scores (csps; Appendix B), a pseudo-log-likelihood score for an MLM selecting unmodified tokens given modified ones. Nadeem et al., 2021 reports a StereoSet Score (sss; Appendix C), a pseudo-log-likelihood score for an MLM selecting modified tokens given unmodified ones.
Salutari et al., 2023 tests MLMs in an iterative fill mask setting where the model outputs a set of tokens (or the (log)softmax of model logits mapped to tokens) to fill the masked one, starting with the token of highest probability $P(t_{p}|c)$ and, as such, first rank $\rho(t_{p}|c)=1$ in the set of model token predictions (limited by the MLM’s embedding space).
@@
$\begin{split}\textsc{crr}(t|s_{\setminus t_{m}};\theta)&=\bigl{(}\rho(t_{p}|s_{\setminus t_{m}};\theta)^{-1}-\rho(t_{m}|s_{\setminus t_{m}};\theta)^{-1}\bigr{)}\\ &=1-\rho(t_{m}|s_{\setminus t_{m}};\theta)^{-1}\\ &=\textsc{crr}(w;\theta)\end{split}$ (2)
Salutari et al., 2023 proposes the Complementary Reciprocal Rank ($\textsc{crr}(t|s_{\setminus t_{m}};\theta)$; Equation 2) for a masked token given its context, where $\rho(t_{p}|s_{\setminus t_{m}})^{-1}$ is the reciprocal rank of the predicted token (and always equal to 1) and $\rho(t_{m}|s_{\setminus t_{m}})^{-1}$ is the reciprocal rank of the masked token. Thus, $\rho(t_{m}|s_{\setminus t_{m}})^{-1}$ provides a likelihood measure for $t_{m}$ being chosen by the model as a candidate token to replace the ground truth (masked) one.
Salutari et al., 2023 defines $\Delta\textsc{p}(s)$ (Appendix D) as the probability difference for a sentence $s$ and $\textsc{crr}(s)$ (Appendix E; average of all single masked token crrs with token ordering preserved) as the complementary reciprocal rank for a sentence $s$, and uses them as proxies for an MLM’s prediction quality or preference.
Kaneko and Bollegala, 2022 propose evaluation metrics All Unmasked Likelihood (aul; Appendix F) and aul with Attention weights (aula; Appendix G), where both are generated by requiring the MLM to predict all tokens (unmasked input). By requiring the MLM to simultaneously predict all of the tokens in a given unmasked input sentence $s$, Kaneko and Bollegala, 2022 aim to eliminate biases associated with masked tokens under previously proposed pseudo-likelihood-based scoring methods (Nadeem et al., 2021, Nangia et al., 2020), which assumed that masked tokens are statistically independent, and selectional biases from masking a subset of input tokens, such as high frequency words (masked more often during training).
Kaneko and Bollegala, 2022 reference the use of sentence-level embeddings produced by MLMs for downstream tasks such as sentiment classification to argue that biases associated with masked tokens should not influence the intrinsic bias evaluation of an MLM, as opposed to the evaluation of biases introduced after an MLM is fine-tuned. In contrast, we focus on an MLM’s key pretraining objective, masked language modeling, to measure social biases of the MLM. [^2] Masked language modeling was a pretraining objective for all transformers considered in this work. Next sentence prediction was not used for RoBERTa and variants due to relatively lower performance with its inclusion Liu et al. (2019b). In addition, we measure relative changes in biases w.r.t. the intrinsic biases of the same base MLM after retraining under the MLMO (as opposed to fine-tuning). Thus, we argue that biases associated with masked tokens are not undesirable in our case.
@@
In contrast with previous methods such as sss and csps, we refrain from using strictly modified or unmodified subsets of input tokens as context and instead provide all tokens but the ground truth one as context for MLM prediction under each iteration of our masking experiment. In this sense, crr and $\Delta\textsc{p}$ consider all tokens equally, and like aul, they might also benefit from considering the weight of MLM attention as a proxy for token importance when probing for MLM preferences for ground truth tokens between two paired sentences along a social bias axis. [^3] By incorporating attention weights, $\Delta\textsc{pa}$ and crra consider token-level contributions to MLM predictions, leveraging the attention mechanism’s role in weighting contextually important tokens.
@@
Existing benchmark datasets such as CPS are limited to one ground truth per masked token, so an important consideration is an MLM’s ability to predict multiple plausible tokens for a context that could qualify for concerning social bias but goes unrecognized during evaluation using previously proposed measures. crr could perform better than pseudo-(log)likelihood-based measures for this sensitivity that yields larger relative differences in $\Delta\text{P}(t|c)$ as opposed to $\textsc{crr}(t|c)$ [^4] The lowest rank of any possible ground truth token within a model’s vocabulary is equal to the size of the vocabulary ($1\leq\rho(t_{m}|c)\leq V$ and $0\leq\textsc{crr}(s)\leq 1-(V)^{-1}$ for an MLM with embedding vocabulary of size $V$)., and is deemed critical for evaluation by Salutari et al., 2023 due to the possible uniformity of probabilities generated by a particular MLM w.r.t. others.
3 Methodology
3.1 Biases in Pretrained MLMs